# AI-Powered Episodic Future Thinking

### v1

@@ 3.1 Participants
Following approval from the Institutional Review Board (IRB), the EFT research team carried out an evaluation study with 25 participants, specifically targeting individuals with obesity. Eligibility was assessed based on the following criteria: a BMI of 30 or higher, fluency in English, and an age range of 18 to 65 years. Participants were invited by email from a database of individuals who had previously completed eligibility questionnaires for studies at the center for our prior studies. All participants had provided consent to be contacted about research studies. The recruitment process began with 346 email invitations of which 17.9% (62 individuals) started the screening questionnaire—to evaluate eligibility. Out of those, 15.3% (53 individuals) completed the questionnaire, resulting in 9.8% (34 individuals) being deemed eligible and providing consent to participate. Ultimately, 7.2% (25 individuals) completed the study. The demographics of the participants are detailed in Table 1. Household size is summarized using the median and interquartile range (IQR), where the median represents the middle value and the IQR (25th–75th percentile) shows the range in which the central 50% of the data falls. In general, the set of participants seemed to be relatively balanced across key demographic characteristics.
@@
Characteristic Complete (n=25)
BMI (kg/m2), mean (SD) 38.8 (8.1)
Age, mean (SD) 38.5 (10.1)
Gender, n (%)
Male 10 (38.5)
Female 15 (61.5)
Household income, n (%)
<<<50,000 7 (30.1)
\$50,000–\$79,999 6 (23.1)
\$80,000–\$119,999 7 (27.0)
$\geq$120,000 5 (19.2)
Household size, median (IQR) 2 (2, 4)
Education, n (%)
HS/GED or less 7 (30.1)
Some college 6 (23.1)
2-yr college degree 3 (11.5)
4-yr college degree 7 (26.9)
Grad/professional degree 2 (7.7)
Race, n (%)
White 17 (65.4)
Black/AA 3 (15.4)
Asian 1 (3.8)
More than one race 4 (15.4)
Ethnicity, n (%)
Hispanic/Latino 2 (7.7)
@@
Users began by completing an informed consent process on the Qualtrics platform [61] where they provided their demographic information. They were then redirected to our website, where each user was assigned a unique ID stored in the database for use during their interactions.
@@ 3.2 Data Collection
In this study, one of the goals was to better understand people’s experiences with EFTeacher, particularly their interactions with it. To do this, we used thematic analysis to perform a qualitative analysis, identifying themes from open-ended survey responses. This approach was inductive, meaning themes were drawn directly from the participants’ responses without being predetermined.
@@
After these ratings, participants were asked to elaborate on why they felt the way they did. They were also asked to identify features of EFTeacher they found helpful and engaging, as well as those they found frustrating or less useful. This included listing at least two features they liked and believed should remain in future versions of the chatbot and two features they disliked and thought should be improved. Finally, participants could share any additional feedback or thoughts about their overall experience with EFTeacher. This process aimed to gather detailed insights into the user experience, including both positive and negative aspects. This feedback would help refine EFTeacher and inform the design and evaluation of similar chatbots in the future. Each participant is assigned an ID number from 100-150.
Quantitative data was also collected for both SUS and content characteristics. For the SUS [29], we calculated the mean and standard deviation (SD) across all 25 users. For key content characteristics (i.e., Vividness, Liking, Excitement, and Importance), we calculated the median of these four items from each EFT cue that was generated for a specific timeframe, where users rated each item on a scale of 1 (not at all) to 5 (extremely) based on how they felt about each one.
3.3 Users’ Overall Experience
To address RQ1 (i.e., What are users’ overall experiences of using EFTeacher?), the usability of the chatbot was assessed using 10 standard SUS items on a 5-point Likert scale. Figure 2 illustrates the distribution of SUS scores. The average SUS average score is 89.93. with a standard deviation of approximately 10.01, indicating that participants generally provided high and consistent usability ratings. Previous research has examined the interpretation of SUS scores in usability assessments, highlighting that higher scores are associated with superior usability experiences [62]. Bangor et al. suggest that products with scores above 70 are generally considered usable, while those in the high 70s to upper 80s indicate better usability. Truly superior products tend to score above 90, reflecting strong user satisfaction. The chatbot’s score of 89.93 places it well within this range, suggesting a highly positive user experience.
@@ 3.4 Assessment Regarding Key Content Characteristics
To address RQ2 (i.e., What are users impressions of using EFTeacher?), we focused on four key characteristics of the content of the cue texts. Figure 3 displays four violin plots representing ratings over different time frames (1 month, 3 months, 6 months, 1 year, 3 years, 5 years, and 10 years) for the characteristics of Vividness, Liking, Excitement, and Importance. Each plot shows the distribution of ratings on a scale from 1 to 5, with the width of each violin indicating the density of responses at each rating level.
Ratings for vividness are consistently high across all time frames, with most responses clustered around 4 to 5. This pattern indicates a strong and enduring perception of vividness, with a median rating close to 5 across all intervals.
Similarly, ratings for liking  [^1] How much the participant likes the event show high consistency, predominantly centered near 5 across all time frames. This distribution suggests a lasting sense of appreciation or a positive preference, with little variation over time.
The excitement plot exhibits slightly more variability than was in evidence for vividness and liking. For shorter time frames, such as 1 month and 3 months, ratings are more dispersed, ranging between 3 and 5. However, as time progresses, ratings consolidate towards higher values, with most responses centered around 4 to 5 at longer intervals (e.g., 10 years), indicating increasing consistency and sustained excitement over time.
The importance plot shows a slightly wider spread of ratings for shorter time frames, particularly for 1 month and 3 months, where responses are more varied. Over longer durations, ratings converge near 4 to 5, reflecting a growing or stable recognition of importance with time.
The scores users assigned to cue texts regarding the four key content characteristics highlights the consistently high median ratings for vividness, liking, excitement, and importance across all timeframes, demonstrating that users perceive EFTeacher as impactful based on what it generated for them through their interactions with it.
Figure 3: Ratings for vividness, liking, excitement, and importance characteristics.
@@ 3.5 Assessment Regarding User Perceptions
To address RQ3 (i.e., How do users perceive working with an interactive AI chatbot for EFT generation?), the research team employed thematic analysis, i.e., a systematic approach to identify themes from user feedback [32, 63, 64]. Three researchers, each with different areas of expertise, independently reviewed the participants’ open-ended responses multiple times. They each developed a list of codes, which are short labels summarizing key points from the feedback. A facilitator, who had expertise in qualitative methods but limited prior knowledge of the survey (to reduce bias), gathered these codes and compiled them into a comprehensive codebook.
The facilitator and coders then held collaborative sessions to refine the codes, merge similar ones, and identify recurring patterns. This iterative process led to the formation of broader thematic categories. For each theme, the team defined a description and selected representative quotes from user responses to illustrate it. Ultimately, five key categories emerged, which were discussed extensively and finalized in a cohesive codebook.
@@
Using this finalized codebook, the coders reviewed the user responses again, selecting excerpts that best exemplified each theme. The facilitator aggregated these responses and calculated the reliability between coders, a measure of how consistently the coders applied the codes, which was found to be very high, i.e., 94%. The team then agreed on a final table of themes, including example quotes to support each one.
@@
The analysis provided insights into participants’ experiences with EFTeacher, focusing on: communication, adaptability, imagination, reflection, and its interface. These findings highlighted varying levels of engagement and interaction, offering valuable feedback for understanding how users perceived and interacted with the chatbot. The next five subsections explain these themes, providing a framework for understanding participant experiences and interactions with EFTeacher throughout the study. To provide details, quotes from participants are included, with the participant number given for attribution purposes.
3.5.1 Communication
This theme focuses on descriptions of dialogue and communication with EFTeacher. The participants decribe how conversational or human-like their interactions with EFTeacher felt. They widely praised EFTeacher for its ability to engage in human-like conversations. Its tone was frequently described as natural, and exceeding user expectations. Many noted that the dialogue felt impressively lifelike and engaging:
@@ 3.5.1 Communication
“Actually, the response was better than I would expect a human to respond.” [P105]
“The words that EFTeacher used were just everyday words, nothing fancy or unnatural. This didn’t sound like it was coming from a ’machine’; to me, it sounded natural. I felt as if I was having a conversation with an instructor!” [P112]
Participants emphasized the model’s ability to maintain a conversational flow that mirrored human interactions, often feeling natural [P139], genuine [P137], and incorporating leading questions [P117]. The responses were also praised for being empathetic and relevant:
“It really interacted with me like a human! I expressed so much joy at something it wrote and said I might need to save it to read later, and it responded so thoughtfully.” [P118]
“Frankly, given the same questions (somewhat rote), a smart human being with the same objective as the AI obviously had would come up with similar text.” [P103]
Despite its strengths, some participants found the interactions occasionally too polished or scripted, detracting from the perceived naturalness:
“I think the responses are arguably too well-written to reflect the way the average person speaks. I liked the way it worked a lot, but there are some aspects that don’t feel natural.” [P106]
@@
“Even though it was formed from my prompts, it still sounded very scripted.” [P104]
Repetition also emerged as a common critique. While many users appreciated the flow of the conversation, some noted that EFTeacher tended to repeat similar requests for details, making the interaction feel less dynamic:
“Not only was their storytelling great, but it seemed to flow somewhat like a real conversation, except for the repetitiveness of asking for the details it needed each time.” [P123]
“It was very repetitious. I’m not sure a human would respond the exact same way to every question.” [P119]
These insights suggest that while EFTeacher demonstrates a remarkable ability to create engaging and human-like conversations, there is room for improvement. Participants’ feedback highlights a dual expectation: While many users appreciated EFTeacher’s thoughtful and engaging dialogue, the repetition of similar requests for details caused frustration for some. Striking a balance between maintaining high-quality responses and reducing repetitive patterns will be crucial for future iterative refinement of EFTeacher.
3.5.2 Adaptability, including personalization, customization, and flexibility
Participants widely praised EFTeacher for its ability to respond to personalized prompts and tailor outputs to individual preferences. “I like you can change the cue because if there’s more you want to add or take out you can.” [P122]
This flexibility gave users a sense of control over their interactions and enhanced their overall experience. Many found it easy to adjust the tool’s outputs to align with their specific needs:
“It was very easy to have EFTeacher make changes. I didn’t know if I could ask for specific words to be changed, so I just tried it, and it worked! When I asked for less vivid wording, it changed it to something that I liked better.” [P112]
“I enjoyed how individual words could be replaced. I didn’t like an adjective that was used and asked it to replace that specific word. I also liked how the response could be completely rewritten. It’s great that you can scrap a response entirely and have the bot create a new one.” [P116]
Participant [P118] appreciated how EFTeacher adapted to feedback, allowing them to refine outputs in real time. Simple edits such as rearranging phrases or adjusting tone were implemented with ease:
“The word choice it used was kind of strange, (in one particular instance it said that a table was “laden” with food, and it felt very Middle Ages). I brought this up, however, and EFTeacher easily fixed it! I realize that EFTeacher is able to use the full dictionary to describe things - but that doesn’t always mean the words it chooses are commonly used in our modern vernacular. Again, easy to fix though!” [P118]
Participant [P115] highlighted EFTeacher’s ability to paraphrase and confirm their intent, creating a more engaging and collaborative interaction:
“I liked the way it confirmed what you were trying to convey. It was great that EFTeacher understood my words and paraphrased it back to me.” [P115]
However, one of the participants identified limitations in EFTeacher’s ability to fully simplify responses when requested. While the tool was adaptable, certain adjustments fell short of expectations:
“Even when asked to simplify it was still more than I thought a simple description would be. For some people less words work better for their understanding. So when asked to simplify something, I would expect a shorter answer with right-to-the-point statements.” [P125]
Another participant noted EFTeacher’s capacity to enhance narratives without losing the original intent:
“It picked up on me accidentally submitting an answer before finishing my thoughts. I accidentally hit enter instead of deleting a word and it picked up on that and prompted me to continue my thought” [P122]
Most of the feedback underscores the importance of personalization and adaptability in enhancing user interaction, emphasizing the need to meet diverse user expectations.
3.5.3 Imagination
Participants discussed EFTeacher’s facilitation of imagination, e.g., ability to guide imagery and visualization of future events, describing it as enhancing their milestones and making them feel even more meaningful [P120], painting a vivid picture [P125], and triggering their imagination to visualize these moments [P122]. One participant emphasized its impact:
“I was really struggling to imagine events ten years from now, and it helped me string together some of my piecemeal thoughts!” [P118]
Many users praised its descriptive capabilities, noting that the tool effectively transformed abstract ideas into detailed visuals.
“I enjoyed the details added. I am not a creative writing person and it was nice to see things more vividly.” [P109]
“I liked that the extra details provided a larger picture.” [P138]
“It added vivid details that made it much easier to imagine my future” [P106]
“It was very well done, and I was impressed by the small details it would come up with, and the vividness of its language.” [P123]
Participants appreciated how EFTeacher’s thoughtful use of adjectives and descriptive language made their milestones feel more meaningful and polished:
“Used lots of adjectives. The statement by the AI describes things better than I could put it. It helped me imagine actually being there. By using more descriptive wording I felt as if I could really be doing these things and feeling that way.” [P125]
“The flowery words used helped see how great the future may be.” [P115]
Several participants highlighted that the guided prompts and questions posed by EFTeacher were particularly effective in facilitating the imagination of their stories. The questions were described as helpful [P109], well-designed [P103], thought-provoking [P119], and structured yet flexible, allowing for creativity:
@@ 3.5.3 Imagination
“The program/EFTeacher asked perfect questions to help me envision these events!” [P137]
“I liked how it asked questions about my storyline because it made me think about the setting, people, and everything. It really helped bring out the writer in me.” [P122]
Participants also found the tool easy to use, noting that it naturally guided them into crafting imaginative stories. As one participant described:
“It was easy to use, and coax’s it into imaginative stories. It was just very simple, as long as you understood what it was asking for, to use and make it come up with very nice stories.” [P123]
One participant mentioned expecting the chatbot to take the lead in storytelling or provide meaningful new ideas beyond their initial input:
“It felt like I was the one coming up with every story, and they were just riffing on what I came up with by myself. I don’t think this is bad, but I didn’t feel like it was giving me anything to go off of besides the timeframe the story took place in.” [P123]
EFTeacher’s ability to stimulate the imagination through vivid descriptions, questions, and prompts is a standout feature. It has the potential to further enhance its role as a facilitator of imagination and creativity, making it an even more effective tool for users seeking to envision and articulate their future.
3.5.4 Reflection
Some participants, reflecting on their experience, expressed emotions, connections, and results of the imagery elicited by EFTeacher. They shared their key takeaways from interacting with EFTeacher, highlighting how it prompted them to reflect on their goals and aspirations in new ways.
One participant reflected on how EFTeacher helped them think more broadly about their future:
“I don’t typically think very far into the future, and it made me consider where I want my life to be in X months/years.” [P106]
Similarly, another participant noted how EFTeacher prompted thoughts they had not previously explored:
“It made me think about things I really haven’t thought about.” [P122]
“The system made realize what the future may hold.” [P115]
The reflective process elicited by EFTeacher was described as both emotionally impactful and validating. For instance, one participant shared how the experience left them feeling hopeful:
“It made me feel hopeful about these future events. It was also validating some of the feelings that I had.” [P107]
This sentiment was echoed by another participant, who highlighted the excitement generated by envisioning future events:
“ it made you excited for the event.” [P138]
[P107]’s feedback highlighted the vividness of the experience, which played a key role in inspiring hope:
“It was so vivid that it gave me hope.” [P107]
“The AI was on point and gave me a lot of hope that I went to work the next day and could imagine these events” [P107]
EFTeacher’s interactive and specific prompts encouraged self-reflection by helping participants gain meaningful insights from their experiences. Participants mentioned how this approach encouraged self-reflection:
@@ 3.5.4 Reflection
“Being prompted to imagine specific details made me reflect on my life.” [P137]
“It allowed me to express my dreams of the future.” [P127]
“The way the system took my thoughts and made them into a great story. The use of descriptive words brought my true-to-heart feelings to life with little effort on my part.” [P115]
Another participant [P118] reflected on their overall experience with EFTeacher:
“I expressed so much joy at something it wrote… and it responded thoughtfully.” [P118]
One participant appreciated the experience but found that redundancy slightly reduced their enjoyment:
@@
“I really liked this experience. The redundancy made sense, but it was the only aspect that made it feel slightly less enjoyable for me. That said, I understand that not everyone envisions their personal journey in the same way or for the same duration.” [P106]
For another participant, the level of detail felt overwhelming at times:
“Some of the responses contained so much detail that it became overwhelming. There was just a lot of information to process.” [P138]
These observations illustrate how EFTeacher could guide participants to envision their futures with greater clarity. By prompting users to focus on specific details and engaging with their inputs thoughtfully, EFTeacher inspired reflections and a connection to their aspirations.
3.5.5 Interface
Participants shared some insights on EFTeacher’s interface, including its usability and visual design, discussing both its strengths and areas for improvement. While many found the platform intuitive and easy to navigate, others highlighted specific challenges that impacted usability and accessibility.
One common suggestion was to improve the design for better readability, and suggestions for better color contrast were common. For example, participant [P129] commented on the color scheme and suggested improved audio as well as visual support:
“The color scheme needed improvement, as some words blended into the background, making them hard to see.” [P129]
“no option for audio and visual. Hearing the descriptions and scenarios aloud would have added good detail” [P129]
Beyond readability, participants proposed features to enhance the interface’s usability. For instance, the inclusion of a word count feature was suggested to help users meet expectations for detail:
“If there is an expectation of level of detail or goal, then it would be nice to track that from the start.” [P117]
Another participant expressed difficulty with copying text, emphasizing the need for smoother interaction:

### v2

@@ 3.1 Participants
Characteristic Complete (n=25) Characteristic Complete (n=25)
BMI (kg/m2), mean (SD) 38.8 (8.1) Education, n (%)
Age, mean (SD) 38.5 (10.1) HS/GED or less 7 (30.1)
Gender, n (%) Some college 6 (23.1)
Male 10 (38.5) 2-yr college degree 3 (11.5)
Female 15 (61.5) 4-yr college degree 7 (26.9)
Household income, n (%) Grad/professional degree 2 (7.7)
<<<50,000 7 (30.1) Race, n (%)
\$50,000–\$79,999 6 (23.1) White 17 (65.4)
\$80,000–\$119,999 7 (27.0) Black/AA 3 (15.4)
$\geq$120,000 5 (19.2) Asian 1 (3.8)
Household size, median (IQR) 2 (2, 4) More than one race 4 (15.4)
Ethnicity, n (%) Hispanic/Latino 2 (7.7)
@@
Following approval from the Institutional Review Board (IRB), and extensive pilot testing, the EFT research team carried out an evaluation study with 25 participants, specifically targeting individuals with obesity. Eligibility was assessed based on the following inclusion criteria: a BMI (measured in kg/m) of 30 or higher, fluency in English, and an age range of 18 to 65 years. Participants were invited by email from a database of individuals who had previously completed eligibility questionnaires for our many prior studies. All participants had provided consent to be contacted about research studies. The recruitment process began with 346 email invitations of which 17.9% (62 individuals) started the screening questionnaire—to evaluate eligibility. Out of those, 15.3% (53 individuals) completed the questionnaire, resulting in 9.8% (34 individuals) being deemed eligible and providing consent to participate. Ultimately, 7.2% (25 individuals) completed the study. The demographics of the participants are detailed in Table 1. Household size is summarized using the median and interquartile range (IQR), where the median represents the middle value and the IQR (25th–75th percentile) shows the range in which the central 50% of the data falls. In general, the set of participants seemed to be relatively balanced across key demographic characteristics, in part because the recruitment proceeded in small steps, with a shifting focus to help achieve balance.
Users began by completing an informed consent process on the Qualtrics platform [qualtrics] where they provided their demographic information. They were then redirected to our website, where each user was assigned a unique ID stored in the database for use during their interactions.
@@ 3.2 Data Collection
In this study, one of the goals was to better understand people’s experiences with EFTeacher, particularly their interactions with it. To do this, we used thematic analysis to perform a qualitative analysis, identifying themes from open-ended survey responses.
@@
After these ratings, participants were asked to elaborate on why they felt the way they did. They were also asked to identify features of EFTeacher they found helpful and engaging, as well as those they found frustrating or less useful. This included listing at least two features they liked and believed should remain in future versions of the chatbot and two features they disliked and thought should be improved. Finally, participants could share any additional feedback or thoughts about their overall experience with EFTeacher. This process aimed to gather detailed insights into the user experience, including both positive and negative aspects. This feedback would help refine EFTeacher and inform the design and evaluation of similar chatbots in the future. Each participant was assigned an ID number from 100-150.
Data was also collected for both SUS and content characteristics. For the SUS [brooke1996sus], we calculated the mean and standard deviation (SD) across all 25 users. For key content characteristics (i.e., Vividness, Liking, Excitement, and Importance), we calculated the median of these four items from each EFT cue that was generated for a specific timeframe, where users rated each item on a scale of 1 (not at all) to 5 (extremely) based on how they felt about each cue text.
3.3 Users’ Usability Experience
To address RQ1 (i.e., How do users evaluate the usability of EFTeacher?), the usability of the chatbot was assessed using 10 standard SUS items on a 5-point Likert scale. Figure 3 illustrates the distribution of SUS scores. The average SUS average score is 89.93, with a standard deviation of approximately 10.01, indicating that participants generally provided high and consistent usability ratings.
Previous research has examined the interpretation of SUS scores in usability assessments, highlighting that higher scores are associated with superior usability experiences [bangor2008empirical]. Bangor et al. suggest that products with scores above 70 are generally considered usable, while those in the high 70s to upper 80s indicate better usability. Truly superior products tend to score above 90, reflecting strong user satisfaction. The chatbot’s score of 89.93 places it within this range, suggesting a highly positive user experience.
@@ 3.4 Assessment Regarding Key Content Characteristics
To address RQ2 (i.e., What are users’ impressions of the EFT cue texts generated by EFTeacher?), we focused on four key characteristics of the content of the cue texts. Figure 4 displays four frequency distributions (rating values on the x-axis, frequency on the y-axis) for each characteristic, representing ratings over different time frames (1 month, 3 months, 6 months, 1 year, 3 years, 5 years, and 10 years) for the characteristics of Vividness, Liking, Excitement, and Importance. Each plot shows the distribution of ratings on a scale from 1 to 5.
Figure 4: Ratings for vividness, liking, excitement, and importance characteristics
Ratings for Vividness show a strong concentration at the highest rating (5) across all time frames. The frequency bars for rating 5 dominate each time interval, indicating a consistently high perception of vividness. Lower ratings (3 and 4) appear infrequently, supporting the conclusion that participants found the imagined future events vivid.
Ratings for Liking [^1] How much the participant likes the event are also strongly skewed toward 5 across time frames. A modest number of responses fall in the 3–4 range, particularly at shorter intervals (e.g., 1 month and 3 months). Nonetheless, since most ratings are 5, it shows that participants generally found the events likable throughout all timeframes.
For Excitement, while rating 5 remains the most frequent, there is a noticeable presence of lower ratings (especially 2, 3, and 4) at earlier time frames, such as 1 and 3 months. As time frames lengthen (e.g., 5 and 10 years), the distribution becomes more concentrated around 5, implying that participants may find longer-term future events more exciting.
Ratings for Importance are generally high, with most bars clustering at rating 5. However, compared to vividness and liking, this category shows more spread in the 2–4 range, particularly in earlier intervals. Their ratings became higher in the later time frames, indicating a growing sense of importance.
The scores of users assigned to cue texts with respect to the four key content characteristics highlight the consistently high median ratings of vividness, liking, excitement, and importance across all timeframes, demonstrating that users perceive EFTeacher as impactful based on what it generated for them through their interactions with it.
@@ 3.5 Assessment Regarding User Perceptions
To address RQ3 (i.e., How do users perceive working with an interactive AI chatbot for EFT generation?), the research team employed thematic analysis, i.e., a systematic approach to identify themes from user feedback [bowman2023using, coates2021practical, bradley2007qualitative]. Thematic analysis is a qualitative research approach used to identify and interpret patterns or emerging themes within qualitative data, such as open-ended survey responses or interview transcripts [braun2021thematic]. It can be conducted either deductively, drawing on existing theories and prior research, or inductively, allowing themes to emerge directly from the data itself [braun2021thematic]. Thematic analysis is a process that requires researchers to work with the data through repeated review to develop, refine, and validate the identified emerging themes or patterns. This method has become widely used in HCI research to evaluate user experiences, inform iterative design, and interpret qualitative feedback from system interactions. By applying thematic analysis, researchers can gain rich insights into user experience, which is essential for developing user-centered technologies, including conversational agents.
@@
In our study, we applied this approach to analyze participants’ open-ended feedback. Three researchers, each with different areas of expertise, independently reviewed the participants’ open-ended responses multiple times. They each developed a list of codes, which are short labels summarizing key points from the feedback. A facilitator, who had expertise in qualitative methods but limited prior knowledge of the survey (to reduce bias), gathered these codes and compiled them into a comprehensive codebook.
The facilitator and coders then engaged in a collaborative session to review and refine the codes. Through some thematic mappings and group discussion, the team merged overlapping codes, clarified distinctions, and identified patterns across responses. This iterative and collaborative process led to the development of broader themes. For each theme, the group defined a clear description and selected illustrative excerpts from participant responses to exemplify its meaning.
@@
Using this finalized codebook, the coders reviewed the user responses again, selecting excerpts that best exemplified each theme. The facilitator aggregated these responses and calculated the reliability between coders, a measure of how consistently the coders applied the codes. Consistency was found to be very high, i.e., 94%. The team then agreed on a final table of themes, including example excerpts to support each one. The analysis provided insights into participants’ experiences with EFTeacher, highlighting five distinct themes.
@@
These findings highlighted varying levels of engagement and interaction, offering valuable feedback for understanding how users perceived and interacted with the chatbot. The next five subsections explain these themes, providing a framework for understanding participant experiences and interactions with EFTeacher throughout the study. To illustrate key points, excerpts from participant responses are included in the text, with participant numbers provided for attribution. As the number of relevant responses is extensive, not all can be discussed in detail here. For a broader view, Table 2 presents additional representative quotes across all themes.
Table 2: Example Excerpts for Themes
Theme ID Excerpt from Participant Response
1 112 “I felt as if I was having a conversation with an instructor!”
@@
105 “Actually, the response was better than I would expect a human to respond.”
@@
104 “Even though it was formed from my prompts, it still sounded very scripted"
120 “It was more descriptive than most humans, but with AI, I feel that is what you want.”
119 “not personal. The repetitiveness made it feel like AI, so not personal."
2 118 “It gave me good prompts and questions to help me fill in blanks and details about potential future situations.”
@@
112 “The program/EFTeacher asked perfect questions to help me envision these events!”
106 “It added vivid details that made it much easier to imagine my future.”
138 “I liked that the extra details provided a larger picture.”
115 “The flowery words used helped see how great the future may be.”
3 118 “I liked how when I suggested simple edits, it complied easily. ”
115 “I liked the way it confirmed what you were trying to convey”
116 “I enjoy how individual words can be replaced. I did not like an adjective that the boy used and asked for it to replace that specific word”
120 “Seems more personable. ChatGPT is overused and not adaptable to the person using it in my opinion.”
139 “Very straight forward and easy to change or add details"
@@
4 137 “Being prompted to imagine specific details made me reflect on my life”
120 “It was helpful seeing all my events/milestones laid out in a beautiful way"
@@
106 “really liked this experience. The redundancy made sense, but it was the only aspect that made it feel slightly less enjoyable for me. That said, I understand that not everyone envisions their personal journey in the same way or for the same duration. ”
125 “Helped my imagination actually being there. By using more descriptive wording, I felt as if I could really being doing these things and feeling that way"
5 129 “could have a better color scheme; felt like everything kind of blended together with the words at times.”